OpenStudy (vshiroky):

Find an equation of the line that passes through (0,2) and is parallel to y=1/2x+7

OpenStudy (luigi0210):

Parallel meaning the same slope: \[\LARGE m=\frac{1}{2}\] Use the point slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\]